OpenStudy (anonymous):

Please help, please. Fan and Medal Jack wants to find the solution to a system of equations using y = x + 3 and y = x2 + 6x - 21. Betty says that Jack can solve x2 + 5x - 24 = 0 to find the x-coordinates of the solutions of the system. Explain and demonstrate why Betty is correct.

OpenStudy (anonymous):

@izthewiz7 to solve y = x + 3 with y = x^2 + 6x - 21, we put them equal to each other to get x^2 + 6x - 21 = x + 3. putting all the terms on the left gives x^2 + 6x - 21 - x - 3 = 0 which gives x^2 + 5x - 24 = 0. so betty is correct.